OpenStudy (aaronandyson):

the sum of the square of two positive integers is 208.If the square of the larger number is 18 times the smaller one.Find the numbers

OpenStudy (aaronandyson):

@butterflydreamer @mathslover

OpenStudy (sepeario):

\[x^2+y^2=200\]\[x^2=18y^2\] Using simultaneous equations, \[18y^2+y^2=200\]\[19y^2=200\]

OpenStudy (sepeario):

Can you solve from here?

OpenStudy (aaronandyson):

Its 208 and not 200

OpenStudy (xapproachesinfinity):

should be x^2=18y not 18y^2?

OpenStudy (sepeario):

The question is extremely badly written. It is ambiguous in the sense that it doesn't clarify what you just said above.

OpenStudy (sepeario):

The answer should come to approx. 3.3 if you did your calculations right.

OpenStudy (sepeario):

For y.

OpenStudy (xapproachesinfinity):

at any rate you got the concept :)

OpenStudy (aaronandyson):

@amistre64

OpenStudy (amistre64):

ive got nothing new to add

OpenStudy (jhannybean):

the sum of the square of two positive integers is 208.If the square of the larger number is `18 times the smaller one`. Find the numbers `smaller one` referring to the smaller of the two integers, either x or y. I think :D Exactly what @xapproachesinfinity stated earlier.